Heather’s grandparents have won two permanent tickets right behind home plate at Yankee Stadium in a recent promotion. Her grandparents intend to go to the games for the next two years and then give the permanent tickets to Heather for the third season and thereafter. If we assume that the retail sale price of the tickets will remain at $15,390 per year (realized at the end of each season) and the discount rate is 5%, what is the value of the inheritance today?
Answer #1

the value of the inheritance today

= the value of the perpetuity

= C / r = 15,390 / 5%

= $  307,800